# Agent Control Protocol ACP v1.17: Admission Control for Agent Actions Technical Specification and Reference Implementation

### 24 Mar 2026 v4

@@ Abstract
Agent Control Protocol (ACP) is a formal technical specification for governance of autonomous agents in B2B institutional environments. ACP is the admission control layer between agent intent and system state mutation: before any agent action reaches execution, it must pass a cryptographic admission check that validates identity, capability scope, delegation chain, and policy compliance simultaneously.
ACP defines the mechanisms of cryptographic identity, capability-based authorization, deterministic risk evaluation, verifiable chained delegation, transitive revocation, and immutable auditing that a system must implement for autonomous agents to operate under explicit institutional control.
ACP operates as an additional layer on top of RBAC and Zero Trust, without replacing them. It is designed specifically for the problem that neither model solves: governing what an autonomous agent can do, under what conditions, with what limits, and with complete traceability for external auditing—including across organizational boundaries.
The v1.17 specification comprises 38 technical documents organized into five conformance levels (L1–L5). It includes a Go reference implementation of 23 packages covering all L1–L4 capabilities, 73 signed conformance test vectors (Ed25519 + SHA-256) plus 65 unsigned RISK-2.0 vectors, an OpenAPI 3.1.0 specification (18 endpoints), an executable payment-agent demonstration (examples/payment-agent/), and an executable multi-organization interoperability demo. v1.17 adds external verifiability: a TLC-runnable TLA+ formal model with three checked invariants, an ACR-1.0 sequence compliance runner (library mode and HTTP mode), five stateful sequence test vectors covering cooldown, F_anom Rule 3, boundary conditions, privilege jump, and benign flow, and an ACP-SIGN-2.0 HYBRID mode stub for the post-quantum transition. It defines 12 explicitly prohibited behaviors whose presence disqualifies conformance at any level, and the mechanisms for interoperability between institutions.
@@ 3  Technical Mechanisms
Each authorization request passes through a deterministic risk function that produces a Risk Score (RS) in the range $[0,100]$. The same input always produces the same result—no stochastic elements, no machine learning in the critical path.
@@ 7.3  Deployment
8  Security Model
ACP explicitly defines which threats it mitigates, which properties it guarantees, and which risks fall outside its scope.
8.1  Threat Model (STRIDE)
@@ 11  Specification and Implementation Status
ACP v1.17 is a complete Draft Standard specification with a full Go reference implementation and external verifiability through a TLC-checked formal model and an ACR-1.0 sequence compliance runner.
11.1  Active Specifications — v1.17 (38 documents)
@@
Table 7: ACP v1.17 active specification documents by conformance level (38 primary documents shown). The remaining 3 documents are the ACP-D decentralized suite (L5, design phase) not yet published. New in v1.16: ACP-RISK-2.0 (L2, deterministic engine v2 with F_anom + Cooldown), ACP-SIGN-2.0 (L1, PQC hybrid spec). New in v1.17: ACR-1.0 sequence compliance runner, TLA+ formal model, ACP-SIGN-2.0 HYBRID stub (pkg/sign2/). Superseded versions (CONF-1.0, CONF-1.1, LEDGER-1.2, REP-1.1, AGENT-SPEC-0.3, DCMA-1.0, CROSS-ORG-1.0, REP-PORTABILITY-1.0) archived in archive/specs/.
@@ 11.3  Conformance Test Vectors — 73 Signed Vectors
Evaluate (APPROVED, no anomaly) 1,012 280 5
Evaluate (DENIED, no anomaly) 863 296 5
Evaluate (all 3 F_anom rules) 1,331 296 5
Evaluate (COOLDOWN short-circuit) 149 112 1
PatternKey (SHA-256 hash) 996 184 4
ShouldEnterCooldown 88 0 0
@@ 12  Evaluation
This section evaluates ACP along three dimensions: (i) computational overhead of the decision function, (ii) behavioral correctness under stateful multi-step scenarios, and (iii) reproducibility and external verifiability of decision outcomes.
@@ 12.2  Experimental Setup
All experiments use the Go reference implementation (v1.17) on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, GOMAXPROCS=8). The evaluation uses three components: the stateless evaluation function (pkg/risk.Evaluate), an in-memory state backend (InMemoryQuerier), and the ACR-1.0 compliance runner (Section 13). Production deployments with persistent storage backends will incur additional I/O latency not measured here.
@@ 12.3  Q1: Computational Overhead
Table 10 reports go test -bench results. The full evaluation path (all three F_anom rules active) completes in 1,331 ns. The cooldown short-circuit path (149 ns) bypasses the risk function entirely.
Answer to Q1: The admission check adds between 149 ns and 1,331 ns of latency per request. Given that the actions ACP governs (API calls, financial transactions, file mutations) operate in the millisecond-to-second range, the overhead is at least three orders of magnitude below the cost of the governed action.
@@ 12.6  Limitations
The evaluation uses an in-memory state backend. Persistent storage, distributed deployments, and concurrent request patterns introduce latency and consistency considerations not measured here. Adversarial robustness under coordinated attack scenarios is not evaluated. These are identified as future work (Section 12 roadmap).
@@ 12.7  Roadmap
Post-quantum Go implementation (Dilithium, circl) v1.18
@@
Table 12: ACP v1.17 development roadmap.
@@ 14  End-to-End Verifiability
ACP v1.17 enables end-to-end verifiability by aligning formal specification, test vectors, and runtime validation through a shared request abstraction. Invariants proven in TLA+ are instantiated as concrete test vectors and validated empirically by the compliance runner.
@@
What is formally verified in TLA+ holds at runtime. Each TLA+ invariant is instantiated as a concrete test vector and validated empirically by the compliance runner. This is the connection across the three layers of the v1.17 verifiability stack: formal $\to$ data $\to$ runtime.
@@ 17  Conclusion
The v1.17 specification is complete. A full Go reference implementation (23 packages, L1–L4), 73 signed + 65 unsigned RISK-2.0 conformance test vectors, 5 stateful sequence test vectors (ACR-1.0 compliance runner), a TLC-runnable TLA+ formal model (3 invariants, 0 violations), an ACP-SIGN-2.0 HYBRID mode stub with Ed25519 and ML-DSA-65 migration path, an executable multi-organization interoperability demo (examples/multi-org-demo/), a payment-agent killer demo (examples/payment-agent/), and an extended OpenAPI 3.1.0 specification (18 endpoints) are publicly available at https://github.com/chelof100/acp-framework-en.
Preprint: https://arxiv.org/abs/2603.18829Zenodo: 10.5281/zenodo.19198298

### 25 Mar 2026 v5

# Agent Control Protocol ACP v1.19: Admission Control for Agent Actions Technical Specification and Reference Implementation

@@ Abstract
[…]
@@ 1.4  The Concrete Scenario ACP Prevents
1.5  Design Insight
ACP is compute-cheap but state-sensitive: decision evaluation is computationally inexpensive, while system performance is primarily constrained by state access patterns.
This separation enables a protocol that is both efficient in isolation ($\sim$820 ns per decision, $\sim$88 ns for the cooldown short-circuit path) and adaptable in deployment: performance can be improved by replacing the state backend without modifying protocol semantics. The LedgerQuerier interface formalizes this boundary—the decision function remains stateless and deterministic; the state backend can be a local in-memory structure for development or a distributed, indexed store in production.
1.6  Contributions
This paper makes the following contributions:
  • Admission control protocol. A deterministic admission control protocol for autonomous agents combining capability, resource, historical, and anomaly-based signals into a single verifiable decision.
  • State-aware execution model. An execution model that explicitly separates decision logic from state management via the LedgerQuerier abstraction, enabling deterministic reasoning about temporal agent behavior.
[…]
  • End-to-end verifiability. A three-layer verifiability pipeline: formal (TLA+, TLC-checked invariants), data (signed conformance test vectors), and runtime (ACR-1.0 compliance runner), enabling independent validation without proprietary infrastructure.
[…]
@@ 3  Technical Mechanisms
Each authorization request passes through a deterministic risk function that produces a Risk Score (RS) in the range $[0,100]$. The same input always produces the same result—no stochastic elements, no machine learning in the critical path. ACP deliberately trades expressiveness for deterministic auditability and reproducibility: every decision can be reconstructed from first principles given only the input request and the current state snapshot.
@@ 7.3  Deployment
8  Threat Model
ACP explicitly defines which threats it mitigates, which properties it guarantees, and which risks fall outside its scope. For the formal adversary model and security properties, see Section 16.
8.1  STRIDE Analysis
@@ 11  Specification and Implementation Status
ACP v1.19 is a complete Draft Standard specification with a full Go reference implementation and external verifiability through a TLC-checked formal model and an ACR-1.0 sequence compliance runner.
11.1  Active Specifications — v1.19 (38 documents)
@@
Table 7: ACP v1.19 active specification documents by conformance level (38 primary documents shown). The remaining 3 documents are the ACP-D decentralized suite (L5, design phase) not yet published. New in v1.16: ACP-RISK-2.0 (L2, deterministic engine v2 with F_anom + Cooldown), ACP-SIGN-2.0 (L1, PQC hybrid spec). New in v1.17: ACR-1.0 sequence compliance runner, TLA+ formal model, ACP-SIGN-2.0 HYBRID stub (pkg/sign2/). New in v1.18: performance benchmarks (latency, throughput, state contention), formal security/threat model, system comparison table. New in v1.19: adversarial evaluation (cooldown evasion, distributed multi-agent attack, state backend stress with Redis backend, compliance/adversarial/). Superseded versions (CONF-1.0, CONF-1.1, LEDGER-1.2, REP-1.1, AGENT-SPEC-0.3, DCMA-1.0, CROSS-ORG-1.0, REP-PORTABILITY-1.0) archived in archive/specs/.
@@ 11.3  Conformance Test Vectors — 73 Signed Vectors
Evaluate (APPROVED, no anomaly) 767 280 5
Evaluate (DENIED, no anomaly) 784 296 5
Evaluate (all 3 F_anom rules) 921 296 5
Evaluate (COOLDOWN short-circuit) 88 112 1
PatternKey (SHA-256 hash) 767 184 4
ShouldEnterCooldown 72 0 0
@@ 12  Evaluation
This section evaluates ACP along four dimensions: (i) computational overhead of the decision function, (ii) behavioral correctness under stateful multi-step scenarios, (iii) reproducibility and external verifiability of decision outcomes, and (iv) robustness of per-agent admission control under adversarial attack patterns.
@@ 12.1  Evaluation Goals
  • Q4: Does ACP maintain admission control effectiveness under adversarial patterns, including cooldown evasion and distributed multi-agent attacks?
@@ 12.2  Experimental Setup
All experiments use the Go reference implementation (v1.19) on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, GOMAXPROCS=8). The evaluation uses three components: the stateless evaluation function (pkg/risk.Evaluate), an in-memory state backend (InMemoryQuerier), and the ACR-1.0 compliance runner (Section 13). Production deployments with persistent storage backends will incur additional I/O latency not measured here.
@@ 12.3  Q1: Computational Overhead
Table 10 reports go test -bench results. The full evaluation path (all three F_anom rules active) completes in 921 ns. The cooldown short-circuit path (88 ns) bypasses the risk function entirely.
Answer to Q1: The admission check adds between 88 ns and 921 ns of latency per request. Given that the actions ACP governs (API calls, financial transactions, file mutations) operate in the millisecond-to-second range, the overhead is at least three orders of magnitude below the cost of the governed action.
12.4  Throughput and State Backend Contention
We evaluate throughput under concurrent load using 10, 100, and 500 simultaneous workers, all sharing a single InMemoryQuerier instance.
Workers ns/op req/s Observation
10 1,087 920,161 Baseline concurrency
100 1,282 780,250 Moderate contention
500 1,404 712,259 High contention
Table 11: ACP throughput under increasing concurrency (Intel Core i7-8665U, Go 1.22, Windows/amd64).
ACP sustains approximately 920,000 requests per second at 10 concurrent workers. Throughput degrades gradually as concurrency increases, reaching approximately 712,000 requests per second at 500 workers. Degradation is gradual rather than catastrophic, and no correctness violations were observed under load.
We further evaluate the impact of agent pool size on throughput (10 workers fixed, AddRequest called per iteration per the execution contract):
Distinct agents ns/op req/s Bottleneck
1 84,791 11,794 O($n$) scan + per-key contention
100 12,533 79,789 State distributed, short lists
1,000 12,628 79,190 Shared mutex ceiling
Table 12: State contention analysis: single-agent workload vs. multi-agent workload.
[…]
These results indicate that the computational cost of the decision function is negligible compared to state access cost. Throughput is primarily bounded by state backend contention, not evaluation complexity. The LedgerQuerier interface is designed to be replaced by production backends (e.g., Redis sorted sets with TTL, time-indexed SQL) that eliminate the O($n$) scan penalty and support horizontal scaling.
@@ 12.6  Q3: Reproducibility and External Verifiability
12.7  Adversarial Evaluation (ACP-RISK-2.0)
We evaluate ACP under the adversary model defined in Section 16. The following attack strategies instantiate specific adversary capabilities, including adaptive request patterns and multi-agent coordination. We conducted three targeted experiments using the same LedgerQuerier abstraction and execution contract as the ACR-1.0 compliance runner, addressing adversarial robustness and real-world state-backend scaling.
Experiment 1: Cooldown Evasion Attack
Setup. A single agent executes 500 requests alternating between high-risk (acp:cap:financial.transfer, restricted; RS =80=80, outcome DENIED) and low-risk (acp:cap:data.read, public; RS =0=0, outcome APPROVED) requests. The adversarial hypothesis is that interleaving approved requests prevents the denial accumulation required to trigger cooldown. This attack instantiates the adversary’s ability to adapt request patterns over time.
Results.
Metric Value
Total requests 500
Approved before cooldown 2
Real DENIED before cooldown 3
Cooldown-blocked 495
Requests processed before first block 5
Throughput 815,927 req/s
Table 14: Experiment 1: Cooldown Evasion Attack (1 agent, 500 requests, InMemoryQuerier, Intel i7-8665U, Go 1.22).
ACP triggered cooldown after exactly 3 real DENIED decisions (requests #0, #2, #4). All subsequent requests—regardless of risk score—were short-circuited to COOLDOWN_ACTIVE at Step 2 of the evaluation pipeline ($\sim$88 ns, see Table 10).
[…]
Experiment 2: Distributed Multi-Agent Attack
Setup. $N\in\{100,500,1\,000\}$ coordinated agents each execute 10 high-risk requests (acp:cap:financial.transfer, restricted; RS =80=80) using distinct agentID values. The adversarial hypothesis is that distributing load across many identities evades per-agent detection thresholds. This attack leverages the adversary’s capability to coordinate multiple agents.
Results.
Agents Total requests Real DENIED Cooldown-blocked Throughput
100 1,000 300 700 478,881 req/s
500 5,000 1,500 3,500 154,111 req/s
1,000 10,000 3,000 7,000 223,376 req/s
Table 15: Experiment 2: Distributed Multi-Agent Attack. Each of the NN agents is blocked after exactly 3 DENIED decisions. Total free denials before full blocking =3N=3N.
Every agent was individually blocked after exactly CooldownTriggerDenials =3=3 real denials.
Interpretation. Per-agent admission control provides strong containment for single-agent threats. For distributed attacks, an attacker with $N$ distinct agent identities can execute $3N$ high-risk requests before all agents are individually blocked. Mitigating this requires cross-agent attribution at the policy layer (e.g., shared-resource rate limiting, identity clustering), which is deliberately outside the scope of ACP’s per-agent admission model. This is a design boundary, not a correctness failure.
Experiment 3: State Backend Stress
Setup. 500 agents execute 20 requests each (10,000 total, fully concurrent) against two backend implementations: the reference InMemoryQuerier (Go mutex over in-process slices) and a minimal RedisQuerier (Redis 7 sorted sets via go-redis/v9, Docker loopback, one Redis command per ledger operation, no pipelining). This scenario exercises the adversary’s ability to generate high request volumes and exploit shared state contention.
Results.
Backend Duration Throughput
InMemoryQuerier $\sim$28 ms $\sim$350,000 req/s
RedisQuerier $\sim$4.6 s $\sim$2,100 req/s
Table 16: Experiment 3: State Backend Stress (500 agents $\times$ 20 requests =10,000=10{,}000 total). InMemoryQuerier variance is high under concurrent load ($\pm$30%) due to Go mutex scheduling; RedisQuerier is stable ($\pm$4%).
Interpretation. The lower throughput of InMemoryQuerier under adversarial load ($\sim$350k req/s) compared to the 500-worker baseline of $\sim$712k req/s (Table 11) is expected: the adversarial workload generates significantly higher contention because every request updates multiple sorted lists under a shared mutex. RedisQuerier eliminates the global Go mutex entirely, trading it for per-operation network RTT. A production-grade Redis backend using command pipelining or server-side Lua scripts would reduce latency to a single round-trip per request. The key architectural result remains: ACP protocol semantics are stable across backends; deployment performance scales with the quality of the state backend, not with the admission control logic itself.
These experiments address the adversarial-robustness and scaling gaps identified in Section 12. All scripts, raw output, and Redis configuration are available in compliance/adversarial/.
Security Properties Under Adversarial Stress
We map the evaluated attack strategies to the security properties defined in Section 16, analyzing whether each property holds under adversarial conditions.
Determinism.
Decision outcomes remain consistent under identical inputs and state conditions, including under concurrent execution and adversarial request patterns.
Cooldown Enforcement.
Attack 1 (cooldown evasion via pattern alternation) directly targets the cooldown mechanism. Results show that cooldown activation cannot be indefinitely avoided: adversarial alternation delays but does not prevent enforcement. This confirms that cooldown is eventually enforced under repeated violations.
Bounded Adaptation Resistance.
Attack 1 also exercises adaptive behavior. The experiment shows that the denial counter evolves monotonically under repeated violations: APPROVED requests do not reset prior denials. As a result, adversarial interleaving strategies cannot asymptotically evade enforcement, establishing a bounded adaptation property.
Per-Agent Isolation.
Attack 2 (distributed multi-agent attack) targets system-wide coordination. Results show that ACP enforces constraints independently per agent, maintaining isolation. However, coordinated behavior across agents is not detected, which constitutes an explicit design limitation rather than a violation.
State Sensitivity Under Load.
Attack 3 (state backend contention) evaluates system behavior under adversarial load. While throughput degrades due to shared state contention, decision correctness and enforcement properties remain unaffected.
Summary. The evaluated attacks confirm that ACP enforces its core security properties under adversarial conditions. No violations are observed; instead, limitations arise from explicit design trade-offs, particularly in cross-agent correlation. These results reinforce the central design insight: ACP is compute-cheap but state-sensitive.
@@ 12.8  Limitations
The evaluation uses an in-memory state backend for latency benchmarks and a single-node Docker Redis instance for backend comparison. Persistent storage, distributed deployments with replication, and wide-area network latency introduce considerations not measured here.
12.9  Summary
[…]
@@ 12.10  Roadmap
Performance benchmarks (latency, throughput, state contention) v1.18 Complete
Adversarial evaluation (cooldown evasion, distributed multi-agent, state backend stress) v1.19 Complete
Post-quantum Go implementation (Dilithium, circl) v1.20
@@
Table 17: ACP v1.19 development roadmap.
@@ 14  End-to-End Verifiability
ACP v1.19 enables end-to-end verifiability by aligning formal specification, test vectors, and runtime validation through a shared request abstraction. Invariants proven in TLA+ are instantiated as concrete test vectors and validated empirically by the compliance runner.
@@
What is formally verified in TLA+ holds at runtime. Each TLA+ invariant is instantiated as a concrete test vector and validated empirically by the compliance runner. This is the connection across the three layers of the v1.19 verifiability stack: formal $\to$ data $\to$ runtime.
@@ 15.3  What ACP Does Not Prescribe
16  Security Model
This section defines the adversary model and the security properties provided by ACP.
16.1  Adversary Capabilities
We consider a probabilistic adversary $\mathcal{A}$ interacting with the system through agent requests. The adversary is assumed to have the following capabilities:
  • Request control: $\mathcal{A}$ can generate arbitrary sequences of requests on behalf of one or more agents, including high-frequency and adversarial patterns.
  • Adaptive behavior: $\mathcal{A}$ can adapt future requests based on past responses (approval/denial outcomes), enabling iterative probing strategies.
  • Multi-agent coordination: $\mathcal{A}$ may coordinate multiple agents to attempt distributed evasion, including alternating identities or cross-agent patterns.
  • Input manipulation: $\mathcal{A}$ can choose all request fields (capability, resource, context) within the bounds of the protocol specification.
16.2  Adversary Limitations
The following components are trusted and cannot be modified by the adversary:
  • Execution contract: the sequence of operations governing state updates (evaluation followed by state transitions) is enforced correctly.
  • State integrity: the underlying state backend (ledger, cooldown state, pattern counters) is append-only and tamper-resistant.
  • Evaluation function: the decision function (ACP-RISK-2.0) is deterministic and correctly implemented.
  • Cryptographic primitives: signature verification and identity binding are assumed secure under standard assumptions (EUF-CMA security of Ed25519).
This model corresponds to an adversary that can fully control inputs but cannot compromise the integrity of the control system itself.
16.3  Security Properties
Under this model, ACP enforces the following properties:
Determinism.
For a given request and system state, the evaluation function produces a unique and reproducible decision outcome. This prevents adversarial exploitation of non-determinism. The RiskDeterminism invariant in the TLA+ model formally specifies this property.
State Consistency.
All state transitions follow the execution contract, ensuring that historical behavior is consistently reflected in future decisions.
Cooldown Enforcement.
An adversary cannot bypass cooldown mechanisms through repeated or adaptive requests. Once cooldown conditions are met, subsequent requests are deterministically denied within the cooldown window, regardless of request content. This is confirmed by the SEQ-COOLDOWN-001 sequence test vector.
History Integrity.
Past decisions and events cannot be modified or removed. The append-only ledger (ACP-LEDGER-1.3) and the LedgerAppendOnlyTemporal TLA+ invariant formally guarantee monotonic accumulation of historical state.
Bounded Adaptation Resistance.
While $\mathcal{A}$ may adapt request patterns, such adaptations cannot eliminate the influence of accumulated history (repeated denials, pattern counts) on future decisions. In particular, the denial counter evolves monotonically: intervening approved requests do not reset prior denials, ensuring eventual enforcement under repeated violations.
Per-Agent Isolation.
Policy enforcement is applied independently per agent. Actions by one agent do not affect the enforcement state of another. This property bounds the impact of a compromised or malicious agent to its own request history.
16.4  Informal Security Argument
Because ACP separates stateless evaluation from state management and enforces a strict execution contract, every request results in a deterministic state transition that cannot be bypassed. The append-only nature of the state backend ensures that historical information accumulates monotonically: cooldown activation follows deterministic threshold conditions applied to accumulated state; repeated probing eventually triggers enforced denial periods.
Because decisions are reproducible given the same input sequence and initial state, any observed behavior can be externally verified. This limits $\mathcal{A}$’s ability to exploit hidden or undocumented system behavior.
16.5  Limitations
The security guarantees of ACP rely on the integrity of the execution environment:
  • Compromise of the state backend (deletion or modification of history) breaks core guarantees.
  • Incorrect implementation of the execution contract may lead to inconsistent behavior.
  • The model does not address side-channel attacks or network-level adversaries.
  • Post-quantum signature migration (ACP-SIGN-2.0 HYBRID, Ed25519 $\to$ ML-DSA-65) is specified and stubbed; full implementation is deferred to v1.20.
ACP is therefore best understood as a control-layer mechanism that assumes a trusted execution and storage environment, while providing strong guarantees on decision consistency and temporal behavior under adversarial inputs.
@@ 17.2  Runtime Policy Systems and Agent Frameworks
Recent work has begun to address security governance for LLM-based agentic systems. Syros et al. [15] propose SAGA, a security architecture that provides user-controlled access policies and cryptographic mediation for agent-to-agent tool use and inter-agent communication. ACP complements this line of work by focusing on deterministic, auditable admission control at the individual request level: rather than governing inter-agent communication channels, ACP enforces stateful risk evaluation and cooldown constraints on each action request, with an externally verifiable execution contract and formal invariants.
@@ 17.5  Summary
Table 20 summarizes the properties of ACP relative to existing approaches.
Property RBAC ABAC OPA Cedar MCP/A2A ACP
Stateful decisions $\times$ $\times$ $\times$ $\times$ $\times$
Deterministic evaluation $\times$
Auditability partial partial partial partial $\times$
Runtime enforcement partial
Temporal behavior $\times$ $\times$ $\times$ $\times$ $\times$
External verifiability $\times$ $\times$ $\times$ $\times$ $\times$
Execution contract $\times$ $\times$ $\times$ $\times$ $\times$
Agent-native $\times$ $\times$ $\times$ $\times$ partial
Table 20: Comparison of ACP with existing access control and agent governance approaches. ✓ = fully supported; $\times$ = not supported; partial = partially addressed.
@@
Unlike existing systems, ACP explicitly models temporal behavior—history accumulation, anomaly detection, and cooldown enforcement—within the admission decision itself, and provides verifiable execution semantics through signed test vectors and a runtime compliance runner. This combination of temporal state, runtime enforcement, and external verifiability is not jointly addressed by any prior system in Table 20.
@@ 18  Conclusion
The v1.19 specification is complete. A full Go reference implementation (23 packages, L1–L4), 73 signed + 65 unsigned RISK-2.0 conformance test vectors, 5 stateful sequence test vectors (ACR-1.0 compliance runner), a TLC-runnable TLA+ formal model (3 invariants, 0 violations), an ACP-SIGN-2.0 HYBRID mode stub with Ed25519 and ML-DSA-65 migration path, an executable multi-organization interoperability demo (examples/multi-org-demo/), a payment-agent killer demo (examples/payment-agent/), and an extended OpenAPI 3.1.0 specification (18 endpoints) are publicly available at https://github.com/chelof100/acp-framework-en.
Preprint: https://arxiv.org/abs/2603.18829Zenodo: 10.5281/zenodo.19219776
@@
ACP demonstrates that admission control for autonomous agents can be both computationally efficient and operationally enforceable—provided that decision logic and state management are cleanly separated. The protocol is compute-cheap but state-sensitive: the LedgerQuerier abstraction keeps the evaluation function at $\sim$820 ns while allowing the state backend to be replaced, optimized, or scaled independently. This is not an implementation detail—it is the architectural property that makes ACP deployable in production without sacrificing governance correctness.